# Independent sets of non-geometric lattices and the maximal adjoint

### v2

@@ 1. $(1\Rightarrow 2)$
    If $\mathcal{M}^{\triangle}$ is an adjoint of $\mathcal{M}$ then by Lemma [3.5] we have $I((\mathcal{L}^{*})^{opp})\subseteq\mathcal{M}^{\triangle}$. Let $\mathcal{A}\notin\mathcal{S}(\mathcal{M})$ and $I$ be a geometrically independent set of $(\mathcal{L}^{*})^{opp}$ spanning $\cap_{C\in\mathcal{A}}C^{c}$. By Theorem [3.7], we have $\left|I\right|=rank_{\mathcal{M}^{\triangle}}(\overline{\mathcal{A}})<\left|\mathcal{A}\right|$, which means $\mathcal{A}\notin\mathcal{M}^{\triangle}$.
@@ 2. $(2\Rightarrow 3)$
[…]
@@ 3. $(3\Rightarrow 1)$
    As we know geometric lattices are co-atomic and we have $\left\{C\mid i\notin C\in\mathcal{C}\left(\mathcal{M}\right)\right\}$ being the hyperplanes of $(\mathcal{L}^{*})^{opp}$ we must have $(\mathcal{L}^{*})^{opp}$ order embedded onto $L(\mathcal{M}^{\triangle})$. As $A(L(\mathcal{M}^{\triangle}))$ is a partition of $\mathcal{C}(\mathcal{M})$ and $\mathcal{M}^{\triangle}$ is of rank $n-d$ we must have the embedding order preserving and a bijection if restricted to $A((\mathcal{L}^{*})^{opp})$.
@@ 5.2. Corollary:
We can see that if $I((\mathcal{L}^{*})^{opp})$ or $\mathcal{S}(\mathcal{M)}$ are matroids, then $\mathcal{M}$ has an adjoint. Moreover, $I((\mathcal{L}^{*})^{opp})$ will be a minimal adjoint and $\mathcal{S}(\mathcal{M)}$ the maximal adjoint in the weak order on the class of matroids with ground set $\mathcal{C}(\mathcal{M})$.
Two separate works, one by James Oxley and Suijie Wang and the second by Relinde Jurrius and Ruud Pellikaan, define an adjoint for linear matroids. We will use the notation of Oxley and Wang as Jurrius and Pellikaan used a dual definition.
Let $\mathcal{M}$ a representable matroid on the ground set $E=(e_{1},...,e_{n})$, and let $\varphi:E\rightarrow\mathbb{F}^{m}$ be a representation of $\mathcal{M}$. To each circuit $C$ of $\mathcal{M}$ is associated with respect to $\varphi$ a unique vector (up to scalar multiplication) $v_{C}=(c_{1},...,c_{n})\in\mathbb{F}^{m}$ such that $\sum_{i=1}^{n}c_{i}\varphi\left(e_{i}\right)=0$, where $c_{i}\neq 0$ iff $i\in C$.
@@ 5.3. Definition:
We first show that a collection of circuits $\mathcal{A}$ of a linear matroid $\mathcal{M}$ corresponding to an independent set of $I((\mathcal{L}^{*})^{opp})$ is independent in the derived matroid for any representation. As $\mathcal{A}$ corresponds to an independent set, we can order the circuits $(C_{1},...,C_{k})$ such that:
@@
    If $\mathcal{A}$ is a minimal set of circuits not in $\mathcal{S}(\mathcal{M})$ then $\left|\cup_{C\in\mathcal{A}}C\right|=rank_{\mathcal{M}}(\cup_{C\in\mathcal{A}}C)+\left|\mathcal{A}\right|+1$ and for any $C\in\mathcal{A}$ we have $\left|\cup_{C\in\mathcal{A}}C\right|=\left|\cup_{C\in\mathcal{A}\setminus\left\{C^{\prime}\right\}}C\right|$.
@@
    If $C,C^{\prime}$ are two circuits such that $\left|C\cup C^{\prime}\right|=rank_{\mathcal{M}}(C\cup C^{\prime})+1$ then for every $e\in C\cap C^{\prime}$ there exists a unique circuit $C^{\prime\prime}$ such that $C^{\prime\prime}\subseteq C\cup C^{\prime}\setminus\left\{e\right\}$ and $C,C^{\prime},C^{\prime\prime}$ are dependent in $\delta\mathcal{M}$.
We proceed now by induction on the number of circuits in $\mathcal{A}$ with the base case being the second observation. Let $\mathcal{A}$ be independent and a minimal set of circuits not in $\mathcal{S}(\mathcal{M})$, use the second observation to eliminate an element $e\in C\in\mathcal{A}$ by replacing every $e\in C^{\prime}\neq C$ with $C^{\prime\prime}$. We are left with an independent set $\mathcal{A}^{\prime}$ of the same size and rank, as $e\in C\setminus(\cup_{C^{\prime}\in\mathcal{A}^{\prime}\setminus\left\{C\right\}}C)$ we see that $\mathcal{A}^{\prime}\setminus\left\{C\right\}$ contains a minimal set of circuits not in $\mathcal{S}(\mathcal{M})$ and the proof follows by the induction hypothesis.
Another old result we can rediscover is that every rank $n-3$ matroid $\mathcal{M}$ has an adjoint. To see this, let =$\mathcal{M}^{\prime}=\left\{\left\{C_{1},C_{2},C_{3}\right\}\mid\cup_{i=1}^{3}C_{i}=[n]\right\}$. It is easy to check that $\mathcal{M}^{\prime}=\mathcal{S}(\mathcal{M})$ and so is an adjoint.
Our next goal is to find an easier way to compute the requirements for being an adjoint, as computing $I((\mathcal{L}^{*})^{opp})$ can be computationally expensive. We start by observing the fact that the fundamental circuit of any basis of $(\mathcal{L}^{*})^{opp}$ must be a circuit of every adjoint of $\mathcal{M}$ as it does not belong to $\mathcal{S}(\mathcal{M})$. We also observe that since $\mathcal{L}^{*}$ is a geometric lattice we have every basis of $(\mathcal{L}^{*})^{opp}$ of the form $(E\setminus B)^{opp}=\left\{C_{B,i}\mid i\notin B\right\}=\mathcal{C}_{B}$ for a basis $B$ of $\mathcal{M}$. The following gives an explicit form to the "fundamental circuits" of $(\mathcal{L}^{*})^{opp}$:
@@ 5.4. Lemma:
If $e\in C\setminus B$ then as $e\notin\cup_{a\in(E\setminus(B\cup\left\{C\right\}))}C_{B,a}$ we must have $(\mathcal{C}\setminus(\left\{C_{B,e}\right\}))\cup\left\{C\right\}$ independent in $(\mathcal{L}^{*})^{opp}$ and so $C_{B,e}\in\mathcal{C}_{\mathcal{C}_{\mathcal{B}},C}$. On the other hand, let $e\notin C\setminus B$ and assume $C_{B,e}\in\mathcal{C}_{\mathcal{C}_{\mathcal{B}},C}$. We then have $\mathcal{C}_{\mathcal{C}_{\mathcal{B}},C}\setminus\left\{C_{B,e}\right\}$ independent with $e\notin\cup_{C^{\prime}\in(\mathcal{C}_{\mathcal{C}_{\mathcal{B}},C})}C^{\prime}$, consequently $\mathcal{C}_{\mathcal{C}_{\mathcal{B}},C}$ is independent.
We continue with an algorithm for computing maximal independent sets of $(\mathcal{L}^{*})^{opp}$ from its bases. Starting with a basis $\mathcal{C}_{B}$ of $(\mathcal{L}^{*})^{opp}$ we have $rank(\mathcal{C}_{B}\setminus\left\{C_{B,i_{1}}\right\})=n-d-1$ for every $i_{1}\notin B$. Therefore, $(\mathcal{C}_{B}\setminus\left\{C_{B,i_{1}}\right\})\cup\left\{C_{1}\right\}\in I((\mathcal{L}^{*})^{opp})$ for every circuit $C_{1}$ of $\mathcal{M}$ containing $i_{1}$. We can then repeat this process, in the $k$’th iteration replacing $C_{B,i_{k}}$ with $C_{k}$ for $k$ such that $i_{k}\in C_{k}\setminus(B\cup(\cup_{j=1}^{k-1}C_{j}))$ and obtain $(\mathcal{C}_{B}\setminus\left\{C_{B,i_{1}},...,C_{B,i_{k}}\right\})\cup\left\{C_{1},..,C_{k}\right\}\in I((\mathcal{L}^{*})^{opp})$.
@@ 5.5. Lemma:
Let $\mathcal{I}$ be a maximal independent set and $\left\{C_{1},...,C_{n-d}\right\}$ an order on its element for which $\omega_{\mathcal{I}}$ exist. By Theorem [4.2] we have a basis $B$ of $\mathcal{M}$ with $E\setminus B=\left\{a_{1},...,a_{d-n}\right\}$ such that $\cap_{i=0}^{k}C_{i}^{c}=\vee_{i=1}^{n-d-k}a_{i}$ for all $t\in\left\{0,..,d-n\right\}$, with the join of the $a_{i}$’s in $\mathcal{L}^{*}$ and we define $\cap_{i=0}^{0}C_{i}=E$ We must then have $\left\{a_{d-n-k-1},...,a_{d-n}\right\}\in C_{k}$ and $\left\{a_{1},...,a_{k}\right\}\notin C_{k}$ for all $k\in[d-n-1]$. Therefore, the algorithm process starting with $\mathcal{C}_{B}$ and replacing the circuits in the defined order $(a_{d-n},...a_{1})$ with the circuits $\left\{C_{1},...,C_{n-d}\right\}$ will result in $\mathcal{I}$.
@@ 5.6. Theorem:
$\mathcal{M}^{\triangle}$ is an adjoint of $\mathcal{M}$ iff $\mathcal{C}_{B}$ is a basis of $\mathcal{M}^{\triangle}$ for every basis $B$ of $\mathcal{M}$ and $\mathcal{C}_{\mathcal{C}_{B},C}$ is a circuit of $\mathcal{M}^{\triangle}$ for every $C\notin\mathcal{C}_{B}$.
@@
We have already seen that $\mathcal{C}_{B}$ is a basis of $\mathcal{M}^{\triangle}$ to see $\mathcal{C}_{\mathcal{C}_{B},C}$ is a circuit of $\mathcal{M}^{\triangle}$ observe that $\mathcal{C}_{\mathcal{C}_{B},C}\notin\mathcal{S}(\mathcal{M})$ and every proper subset of $\mathcal{C}_{\mathcal{C}_{B},C}$ is independent in $(\mathcal{L}^{*})^{opp}$.
For the other direction we first have $\mathcal{M}^{\triangle}$ of the required rank as $\mathcal{C}_{B}$ is a basis, we will now prove that condition $(3)$ of theorem [5.1] hold. Let $e$ be a non-coloop element of $\mathcal{M}$ and $B$ a basis of $\mathcal{M}$ not containing $e$. We will show $\left\{C\mid e\notin C(\mathcal{M})\right\}=\overline{\mathcal{C}_{B}\setminus\left\{C_{B,e}\right\}}$ and so a hyperplane of $\mathcal{M}^{\triangle}$. If $e\in C\in\mathcal{C}(\mathcal{M})$, then $(\mathcal{C}_{B}\setminus\left\{C_{B,e}\right\})\cup\left\{C\right\}\in I((\mathcal{L}^{*})^{opp})$ and so $C\notin\overline{\mathcal{C}_{B}\setminus\left\{C_{B,e}\right\}}$. If $e\notin C\in\mathcal{C}(\mathcal{M})$, then by Lemma [5.4] $C_{B,e}\notin\mathcal{C}_{\mathcal{C}_{B},C}$ and therefore $C\in\overline{\mathcal{C}_{B}\setminus\left\{C_{B,e}\right\}}$.
@@ 6.3. Definition:
In [6] they did not have the criteria for $\delta\mathcal{M}$ being an adjoint of $\mathcal{M}$. Using Theorem [5.1] we know it is enough for $I((\mathcal{L}^{*})^{opp})$ to be independent in $\delta\mathcal{M}$ as $2^{C(\mathcal{M})}\setminus\mathcal{S}(\mathcal{M})$ is already dependent. In particular, if $D(2^{C(\mathcal{M})}\setminus\mathcal{S}(\mathcal{M}))=\\ \uparrow(2^{C(\mathcal{M})}\setminus\mathcal{S}(\mathcal{M}))$ then $\delta\mathcal{M}$ is an adjoint. Two simple examples follow:

### v3

@@ 1. $(1\Rightarrow 2)$
    If $\mathcal{M}^{\triangle}$ is an adjoint of $\mathcal{M}$ then by Lemma 3.7 we have $I((\mathcal{L}^{*})^{opp})\subseteq\mathcal{M}^{\triangle}$. Let $\mathcal{A}\notin\mathcal{S}(\mathcal{M})$ and $I$ be a geometric independent set of $(\mathcal{L}^{*})^{opp}$ spanning $\cap_{C\in\mathcal{A}}C^{c}$. By Theorem 3.8, we have $\left|I\right|=\rho_{\mathcal{M}^{\triangle}}(\mathcal{A})<\left|\mathcal{A}\right|$, which means $\mathcal{A}\notin I(\mathcal{M}^{\triangle})$.
@@ 2. $(2\Rightarrow 3)$
    If $i\in E$ is not a co-loop, then there exists a basis $i\notin B\in\mathcal{M}$. Examining the set of fundamental circuits $C_{B}\setminus\left\{C_{B,i}\right\}$ of $B$, we obtain an independent set of rank nd1n-d-1 contained in $\left\{C\mid i\notin C\in\mathcal{C}\left(\mathcal{M}\right)\right\}$. It is not of full rank as, by the definition of $I((\mathcal{L}^{*})^{opp})$, it will give us an independent set of rank nd+1n-d+1. By the same reasoning, we must have $\left\{C\mid i\notin C\in\mathcal{C}\left(\mathcal{M}\right)\right\}$ closed.
@@ 3. $(3\Rightarrow 1)$
    Since geometric lattices are co-atomic, the set $\left\{C\mid i\notin C\in\mathcal{C}\left(\mathcal{M}\right)\right\}$ corresponds to the hyperplanes of $(\mathcal{L}^{*})^{opp}$. Consequently, $(\mathcal{L}^{*})^{opp}$ must be order-embedded into $L(\mathcal{M}^{\triangle})$. Furthermore, as $A(L(\mathcal{M}^{\triangle}))$ forms a partition of $\mathcal{C}(\mathcal{M})$ and $\mathcal{M}^{\triangle}$ has rank ndn-d , the embedding preserves order and becomes a bijection when restricted to $A((\mathcal{L}^{*})^{opp})$.
@@ 5.3. Corollary:
It follows that if $I((\mathcal{L}^{*})^{opp})$ or $\mathcal{S}(\mathcal{M)}$ are matroids, then $\mathcal{M}$ has an adjoint. Moreover, $I((\mathcal{L}^{*})^{opp})$ serves as the minimal adjoint, while $\mathcal{S}(\mathcal{M)}$ represents the maximal adjoint within the weak order on the class of matroids with ground set $\mathcal{C}(\mathcal{M})$.
Two separate works, [8] by James Oxley and Suijie Wang and [7] by Relinde Jurrius and Ruud Pellikaan, define an adjoint for linear matroids. We will use the notation of Oxley and Wang as Jurrius and Pellikaan used a dual definition.
Let $\mathcal{M}$ a representable matroid on the ground set $E=(e_{1},...,e_{n})$, and let $\varphi:E\rightarrow\mathbb{F}^{m}$ be a representation of $\mathcal{M}$. To each circuit $C$ of $\mathcal{M}$ is associated with respect to $\varphi$ a unique vector (up to scalar multiplication) $v_{C}=(c_{1},...,c_{n})\in\mathbb{F}^{m}$ such that $\sum_{i=1}^{n}c_{i}\varphi\left(e_{i}\right)=0$, where $c_{i}\neq 0$ if and only if $i\in C$.
@@ 5.4. Definition:
We first show that a collection of circuits $\mathcal{A}=\left\{C_{1},...,C_{k}\right\}$ of a linear matroid $\mathcal{M}$ corresponding to an independent set of $I((\mathcal{L}^{*})^{opp})$ is independent in the derived matroid, for any representation. As $\mathcal{A}$ corresponds to an independent set, we can linearly order its circuit such that:
@@
    If $\mathcal{A}$ is a minimal set of circuits not in $\mathcal{S}(\mathcal{M})$ then $\left|\cup_{C\in\mathcal{A}}C\right|=\rho_{\mathcal{M}}(\cup_{C\in\mathcal{A}}C)+\left|\mathcal{A}\right|+1$. Also, for any $C\in\mathcal{A}$ we have $\left|\cup_{C\in\mathcal{A}}C\right|=\left|\cup_{C\in\mathcal{A}\setminus\left\{C^{\prime}\right\}}C\right|$.
@@
    If $C,C^{\prime}$ are two circuits such that $\left|C\cup C^{\prime}\right|=\rho_{\mathcal{M}}(C\cup C^{\prime})+1$ then for every $e\in C\cap C^{\prime}$ there exists a unique circuit $C^{\prime\prime}$ such that $C^{\prime\prime}\subseteq C\cup C^{\prime}\setminus\left\{e\right\}$ and $C,C^{\prime},C^{\prime\prime}$ are dependent in $\delta\mathcal{M}$.
We proceed by induction on the number of circuits in $\mathcal{A}$ , with the base case established by the second observation. Let $\mathcal{A}$ be independent and minimal set of circuits not contained in $\mathcal{S}(\mathcal{M})$. Using the second observation, we eliminate an element $e\in C\in\mathcal{A}$ by replacing each occurrence of $e\in C^{\prime}\neq C$ with $C^{\prime\prime}$. This process results in an independent set $\mathcal{A}^{\prime}$ of identical size and rank. Since $e\in C\setminus(\cup_{C^{\prime}\in\mathcal{A}^{\prime}\setminus\left\{C\right\}}C)$, we observe that $\mathcal{A}^{\prime}\setminus\left\{C\right\}$ still contains a minimal set of circuits not in $\mathcal{S}(\mathcal{M})$. The proof then follows by the induction hypothesis.
An established result we can revisit is that every co-rank $3$ matroid $\mathcal{M}$ has an adjoint. To see this, consider $\mathcal{M}^{\prime}=\left\{\left\{C_{1},C_{2},C_{3}\right\}\mid\cup_{i=1}^{3}C_{i}=[n]\right\}$. A straightforward verification shows that $\mathcal{M}^{\prime}=\mathcal{S}(\mathcal{M})$, confirming that it serves as an adjoint.
Our next objective is to simplify the computation of adjoint requirements, given that evaluating $I((\mathcal{L}^{*})^{opp})$ can be computationally demanding. To begin, we observe that the fundamental circuit of any basis in $(\mathcal{L}^{*})^{opp}$ must be a circuit in every adjoint of $\mathcal{M}$, as it is not contained in $\mathcal{S}(\mathcal{M})$. Additionally, since $\mathcal{L}^{*}$ is a geometric lattice, every basis of $(\mathcal{L}^{*})^{opp}$ takes the form $(E\setminus B)^{opp}=\left\{C_{B,e}\mid e\notin B\right\}=\mathcal{C}_{B}$ for a basis $B$ of $\mathcal{M}$. The following provides an explicit representation of the "fundamental circuits" of $(\mathcal{L}^{*})^{opp}$:
@@ 5.5. Lemma:
If $e\in C\setminus B$ then as $e\notin C_{B,a}$ for every $a\neq e$ we must have $(\mathcal{C}\setminus(\left\{C_{B,e}\right\}))\cup\left\{C\right\}$ independent in $(\mathcal{L}^{*})^{opp}$ and so $C_{B,e}\in\mathcal{C}_{\mathcal{C}_{\mathcal{B}},C}$. On the other hand, let $e\notin C\setminus B$ and assume by negation $C_{B,e}\in\mathcal{C}_{\mathcal{C}_{\mathcal{B}},C}$. We then have $\mathcal{C}_{\mathcal{C}_{\mathcal{B}},C}\setminus\left\{C_{B,e}\right\}$ independent and $e$ not contained in any of its elements, consequently $\mathcal{C}_{\mathcal{C}_{\mathcal{B}},C}$ is independent, contradicting it being a circuit.
We now present an algorithm for computing maximal independent sets of $(\mathcal{L}^{*})^{opp}$ based on its bases. Beginning with a basis $\mathcal{C}_{B}$ of $(\mathcal{L}^{*})^{opp}$, we have $\rho(\mathcal{C}_{B}\setminus\left\{C_{B,i_{1}}\right\})=n-d-1$ for every $i_{1}\notin B$. Consequently, $(\mathcal{C}_{B}\setminus\left\{C_{B,i_{1}}\right\})\cup\left\{C_{1}\right\}\in I((\mathcal{L}^{*})^{opp})$ for any circuit $C_{1}$ of $\mathcal{M}$ that contains $i_{1}$. This process can be iterated: in the $k$’th step, we replace $C_{B,i_{k}}$ with $C_{k}$, where $i_{k}\in C_{k}\setminus(B\cup(\cup_{j=1}^{k-1}C_{j}))$. As a result, we obtain $(\mathcal{C}_{B}\setminus\left\{C_{B,i_{1}},...,C_{B,i_{k}}\right\})\cup\left\{C_{1},..,C_{k}\right\}\in I((\mathcal{L}^{*})^{opp})$.
@@ 5.6. Lemma:
Let $\mathcal{I}$ be a maximal independent set, and consider an ordered sequence $\left\{C_{1},...,C_{n-d}\right\}$ of its elements for which $\omega_{\mathcal{I}}$ exist. By Theorem 4.7, there is a basis $B$ of $\mathcal{M}$ such that $E\setminus B=\left\{a_{1},...,a_{d-n}\right\}$, satisfying $\cap_{i=0}^{k}C_{i}^{c}=\vee_{i=1}^{n-d-k}a_{i}$ for all $t\in\left\{0,..,d-n\right\}$, where the join of the $a_{i}$’s is taken in $\mathcal{L}^{*}$. Additionally, we define $\cap_{i=0}^{0}C_{i}=E$. It follows that $\left\{a_{d-n-k-1},...,a_{d-n}\right\}\in C_{k}$ and $\left\{a_{1},...,a_{k}\right\}\notin C_{k}$ for all $k\in[d-n-1]$. Consequently, the algorithm initiates with $\mathcal{C}_{B}$ and systematically replaces circuits in the given order $(a_{d-n},...a_{1})$ with the circuits $\left\{C_{1},...,C_{n-d}\right\}$, ultimately constructing $\mathcal{I}$.
@@ 5.7. Theorem:
$\mathcal{M}^{\triangle}$ is an adjoint of $\mathcal{M}$ if and only if $\mathcal{C}_{B}$ forms a basis of $\mathcal{M}^{\triangle}$ for every basis $B$ of $\mathcal{M}$, and $\mathcal{C}_{\mathcal{C}_{B},C}$ forms a circuit of $\mathcal{M}^{\triangle}$ for every $C\notin\mathcal{C}_{B}$.
@@
We have already established that $\mathcal{C}_{B}$ forms a basis of $\mathcal{M}^{\triangle}$. To see $\mathcal{C}_{\mathcal{C}_{B},C}$ forms a circuit of $\mathcal{M}^{\triangle}$, note that $\mathcal{C}_{\mathcal{C}_{B},C}\notin\mathcal{S}(\mathcal{M})$ and every proper subset of $\mathcal{C}_{\mathcal{C}_{B},C}$ remains independent in $(\mathcal{L}^{*})^{opp}$.
Conversely, since $\mathcal{C}_{B}$ is a basis of $\mathcal{M}^{\triangle}$, we ensure that $\mathcal{M}^{\triangle}$ has the required rank. Next, we verify that condition $(3)$ of Theorem 5.2 holds. Let $e$ be a non-coloop element of $\mathcal{M}$ and $B$ a basis of $\mathcal{M}$ that does not contain $e$. We aim to show that $\vee\left\{C\mid e\notin C(\mathcal{M})\right\}=\vee(\mathcal{C}_{B}\setminus\left\{C_{B,e}\right\})$ which then forms a hyperplane of $\mathcal{M}^{\triangle}$. If $e\in C\in\mathcal{C}(\mathcal{M})$, then $(\mathcal{C}_{B}\setminus\left\{C_{B,e}\right\})\cup\left\{C\right\}\in I((\mathcal{L}^{*})^{opp})$ and so $C\notin\vee(\mathcal{C}_{B}\setminus\left\{C_{B,e}\right\})$. If $e\notin C\in\mathcal{C}(\mathcal{M})$, then by Lemma 5.5 $C_{B,e}\notin\mathcal{C}_{\mathcal{C}_{B},C}$ and therefore $C\in\vee(\mathcal{C}_{B}\setminus\left\{C_{B,e}\right\})$.
@@ 6.3. Definition:
In [6] they did not have the criteria for $\delta\mathcal{M}$ being an adjoint of $\mathcal{M}$. Using Theorem 5.2 we know it is enough for $I((\mathcal{L}^{*})^{opp})$ to be independent in $\delta\mathcal{M}$ as $2^{C(\mathcal{M})}\setminus\mathcal{S}(\mathcal{M})$ is already dependent. In particular, if $D(2^{C(\mathcal{M})}\setminus\mathcal{S}(\mathcal{M}))=\\ \uparrow(2^{C(\mathcal{M})}\setminus\mathcal{S}(\mathcal{M}))$ then $\delta\mathcal{M}$ is an adjoint. Two simple examples follow: